# Dalitz Plot Kinematics for a Lorentz-Violating Three-Body Decay

### v1

@@ 3.2 Model with Nonzero $c_{13}$
If instead we keep all terms in $c\mathbf{M}_{4\times 4}$, including the $c_{13}$ term, things grow substantially more complicated, and angular approximations will be needed to get tractable analytical expressions. To ease the number of terms in the integrand for the decay rate, we will eventually use a series approximation can be performed in $\theta$ to compute the angular part of the integral.
With the added angular dependence in the square root terms from the modified dispersion relations, two of the three pions dispersion relation become modified to include first order sinusoidal terms. The analog of (17), before $\theta$-integration but with keeping the additional anisotropy term, is
@@
×$\times\bigg{(}4|\vec{p}\,|^{2}\left\{(2+c)|\vec{p}\,|^{2}m^{2}+2m^{4}+|\vec{k}|^{2}\left[2(1+c)|\vec{p}\,|^{2}+(2+c)m^{2}\right]\sin\theta\right\}$
@@
×$\times\left[|\vec{k}|^{2}+m^{2}-\sqrt{|\vec{k}|^{2}+m^{2}}\sqrt{|\vec{k}|^{2}+|\vec{p}\,|^{2}+m^{2}+2|\vec{p}\,||\vec{k}|\cos\theta}\right]\cos 2\theta$
@@
Because of the $\csc\theta$, the $c$-dependent part of the the above integral clearly diverges when the three daughter pions particles are colinear. This is appears to be indicative of the previously noted fact that there may be nonperturbative dependence on $c$ where logarithms arise in the decay rate $\Gamma$; the $\mathcal{O}(c)$ approximations that we have made are inadequate near the poles at $\theta=0$ and $\pi$, where $\cos n\theta=(-1)^{n}$.
@@
To complete the integration without any angular approximations, we need the to find the poles in the boundary equation; we may do this by using (28) below. After insertion of the angle into (24), even when approximating to $\mathcal{O}(c)$ the integral is extremely long. If the approximation is further made of working only to first or second order in the mass $m$, the integral over the daughter momenta becomes more tractable; however, combined with the $\mathcal{O}(c)$ approximation (even with using the modified upper and lower bounds for the integration) the first-order dependence on $m$ vanishes. Thus, to obtain a non-vanishing modification to the decay rate, the integral over on-shell daughter particle energies must be computed to at least $\mathcal{O}(m^{2})$. The integration over one daughter energy yields a hugely complicated expression but with no explicit Lorentz violation terms remaining at leading order. However, after integration over the second on-shell energy, there is a $c$ dependence, and the full decay rate at this order is
@@ 4 Modified Dalitz Plots
$\sqrt{s}(1+c)$displaystyle-- $\left(\sqrt{|\vec{p}\,|^{2}+m^{2}}+\sqrt{|\vec{k}|^{2}+m^{2}}\right)-\frac{c}{2}\left(\frac{m^{2}}{\sqrt{|\vec{k}|^{2}+m^{2}}}+\frac{m^{2}+2|\vec{p}\,|^{2}}{\sqrt{|\vec{p}\,|^{2}+m^{2}}}\right)$ (27)
[…]
@@
The boundary line is plotted in the $m_{23}^{2}$-$m_{13}^{2}$ plane. Setting the quantity $c$ to zero as in (31) naturally gives the standard outline of the plot. In this case, the elongation of the curve depends only on the available decay energy compared with the common mass of the daughters, $\sqrt{s}/m$. When $c$ is increased to a nonzero value the boundary plot shifts according to (32). This is linked to the fact that under Lorentz transformation, the allowed physical region will be rotated and shifted relative to a frame with no Lorentz violation. Increasing the $c$ term in the modified plot is associated with enlarging the physically allowed parameter region, as well as generally shifting it towards smaller values of $m_{13}^{2}$, $m_{23}^{2}$, and (redundantly) $m_{23}^{2}$.
Figure 2 shows the effect of $c$ on the shape of the Dalitz plot for a decay with $\sqrt{s}/m=4.0$, very close to the value for 3.97 for the physical decay $\eta\rightarrow 3\pi^{0}$. Increasing the strength of the Lorentz violation (in this specific model), shifts the kinematically allowed region towards smaller values of $m_{23}^{2}$ and $m_{13}^{2}$—in a symmetric fashion, because the daughter particles are identical. Precision measurements of this particular decay are already recognized to be important, since the matrix element, calculated using chiral perturbation theory, depends directly on the isospin-violating up-down mass difference, $m_{d}-m_{u}$ [41, 42, 43]. Measurements of this and similar processes continue to be used for state-of-the-art measurements of this important parameter in low-energy flavor physics [44, 45].
@@ 5 Conclusions and Outlook
Future analyses of three-body decays involving Lorentz violation may focus particularly on muon decays, $\mu^{\pm}\rightarrow e^{\pm}+\nu+\bar{\nu}$. Muon physics has proved to be a fruitful source for bounds on coefficients in sectors of the SME that are otherwise hard to constrain. In [46], the authors used muon decay as a tool to test the invariance of the weak interaction under Lorentz transformations. Further motivation for investigating muon properties comes from the phenomenology of the muon anomalous magnetic moment (“$g-2$”) [47, 48] and muonic hydrogen [49], where other puzzling deviations from predictions based on the standard model presently exist. Especially for the the magnetic moment experiment, where the direct observable is the decay rate, a full understanding of the three-body decay is important.
The muon work was part of more general development of the use of three-body $\beta$-decays to explore Lorentz violations in the weak sector [50, 51, 52, 53, 54, 55], with the SME coefficients only appearing in the virtual $W^{\pm}$ propagator, rather than the external particle kinematics. The kinematic effects in weak three-body decays have also been studied [56]; however, most kinematic analyses of Lorentz violation in $\beta$-decays have focused on two-body decays, such as $\pi^{+}\rightarrow\mu^{+}+\nu$; or on Cerenkov-like emission of neutrino-antineutrino pairs, as in the normally forbidden process $p^{+}\rightarrow p^{+}+\nu+\bar{\nu}$, for which the kinematics are (because the neutrinos are nearly massless) essentially the same as those for a Cerenkov process.

### v2

@@ 3.2 Model with Nonzero $c_{13}$
If instead we keep all terms in $c\mathbf{M}_{4\times 4}$, including the $c_{13}$ term, things grow substantially more complicated, and angular approximations will be needed to get tractable analytical expressions. To ease the number of terms in the integrand for the decay rate, we will eventually use a series approximation that can be performed in $\theta$ to compute the angular part of the integral.
With the added angular dependence in the square root terms from the modified dispersion relations, two of the three pions dispersion relation become modified to include first-order sinusoidal terms. The analog of (17), before $\theta$-integration but with keeping the additional anisotropy term, is
@@
$\times\bigg(4|\vec{p}\,|^{2}\left\{(2+c)|\vec{p}\,|^{2}m^{2}+2m^{4}+|\vec{k}|^{2}\left[2(1+c)|\vec{p}\,|^{2}+(2+c)m^{2}\right]\sin\theta\right\}$
@@
$\times\left[|\vec{k}|^{2}+m^{2}-\sqrt{|\vec{k}|^{2}+m^{2}}\sqrt{|\vec{k}|^{2}+|\vec{p}\,|^{2}+m^{2}+2|\vec{p}\,||\vec{k}|\cos\theta}\right]\cos 2\theta$
@@
Because of the $\csc\theta$, the $c$-dependent part of the the above integral clearly diverges when the three daughter pions particles are colinear. This appears to be indicative of the previously noted fact that there may be a nonperturbative dependence on $c$ where logarithms arise in the decay rate $\Gamma$; the $\mathcal{O}(c)$ approximations that we have made are inadequate near the poles at $\theta=0$ and $\pi$, where $\cos n\theta=(-1)^{n}$.
@@
To complete the integration without any angular approximations, we need to find the poles in the boundary equation; we may do this by using (28) below. After insertion of the angle into (24), even when approximating to $\mathcal{O}(c)$ the integral is extremely long. If the approximation is further made of working only to first or second order in the mass $m$, the integral over the daughter momenta becomes more tractable; however, combined with the $\mathcal{O}(c)$ approximation (even with using the modified upper and lower bounds for the integration) the first-order dependence on $m$ vanishes. Thus, to obtain a non-vanishing modification to the decay rate, the integral over on-shell daughter particle energies must be computed to at least $\mathcal{O}(m^{2})$. The integration over one daughter energy yields a hugely complicated expression but with no explicit Lorentz violation terms remaining at leading order. However, after integration over the second on-shell energy, there is a $c$ dependence, and the full decay rate at this order is
@@ 4 Modified Dalitz Plots
$\sqrt{s}(1+c)$ - $\left(\sqrt{|\vec{p}\,|^{2}+m^{2}}+\sqrt{|\vec{k}|^{2}+m^{2}}\right)-\frac{c}{2}\left(\frac{m^{2}}{\sqrt{|\vec{k}|^{2}+m^{2}}}+\frac{m^{2}+2|\vec{p}\,|^{2}}{\sqrt{|\vec{p}\,|^{2}+m^{2}}}\right)$ (27)
$\lessgtr$ $\sqrt{|\vec{k}|^{2}+|\vec{p}\,|^{2}+m^{2}\mp 2|\vec{p}\,||\vec{k}|}+\frac{cm^{2}}{2\sqrt{|\vec{k}|^{2}+|\vec{p}\,|^{2}+m^{2}\mp 2|\vec{p}\,||\vec{k}|}}.$
@@
The boundary line is plotted in the $m_{23}^{2}$-$m_{13}^{2}$ plane. Setting the quantity $c$ to zero as in (31) naturally gives the standard outline of the plot. In this case, the elongation of the curve depends only on the available decay energy compared with the common mass of the daughters, $\sqrt{s}/m$. When $c$ is increased to a nonzero value the boundary plot shifts according to (32).
@@
This is linked to the fact that under Lorentz transformation, the allowed physical region will be rotated and shifted relative to a frame with no Lorentz violation. Increasing the $c$ term in the modified plot is associated with enlarging the physically allowed parameter region, as well as generally shifting it towards smaller values of $m_{13}^{2}$, $m_{23}^{2}$, and (redundantly) $m_{23}^{2}$.
Figure 2 shows the effect of $c$ on the shape of the Dalitz plot for a decay with $\sqrt{s}/m=4.0$, very close to the value of 3.97 for the physical decay $\eta\rightarrow 3\pi^{0}$. Figures 2– 4 show the Dalitz plot boundaries for values of the $c_{\mu\nu}$ coefficients at the $10^{-2}$$10^{-1}$ level. These relatively large values are chosen in part to enhance the visual contrast. However, although there are suggestive indications that they might actually be made tighter [42], direct two-sided bounds on the anisotropic quark-sector coefficients that underlie the $c$ coefficient involved here—coming from deep inelastic scattering [43] and Drell-Yan [44] process data—are only at $10^{-4}$$10^{-1}$ levels [45]. So the study of these kinds of Dalitz plots has the potential to be directly useful even at the levels shown in the figures.
Figure 2 shows that increasing the strength of the Lorentz violation (in this specific model), shifts the kinematically allowed region towards smaller values of $m_{23}^{2}$ and $m_{13}^{2}$—in a symmetric fashion, because the daughter particles are identical. Precision measurements of this particular decay are already recognized to be important, since the matrix element, calculated using chiral perturbation theory, depends directly on the isospin-violating up-down mass difference, $m_{d}-m_{u}$ [46, 47, 48]. Measurements of this and similar processes continue to be used for state-of-the-art measurements of this important parameter in low-energy flavor physics [49, 50].
@@ 5 Conclusions and Outlook
Future analyses of three-body decays involving Lorentz violation may focus particularly on muon decays, $\mu^{\pm}\rightarrow e^{\pm}+\nu+\bar{\nu}$. Muon physics has proved to be a fruitful source for bounds on coefficients in sectors of the SME that are otherwise hard to constrain. In [51], the authors used muon decay as a tool to test the invariance of the weak interaction under Lorentz transformations. Further motivation for investigating muon properties comes from the phenomenology of the muon anomalous magnetic moment (“g2g-2”) [52, 53] and muonic hydrogen [54], where other puzzling deviations from predictions based on the standard model presently exist. Especially for the magnetic moment experiment, where the direct observable is the decay rate, a full understanding of the three-body decay is important.
The muon work was part of a more general development of the use of three-body $\beta$-decays to explore Lorentz violations in the weak sector [55, 56, 57, 58, 59, 60], with the SME coefficients only appearing in the virtual $W^{\pm}$ propagator, rather than the external particle kinematics. The kinematic effects in weak three-body decays have also been studied [61]; however, most kinematic analyses of Lorentz violation in $\beta$-decays have focused on two-body decays, such as $\pi^{+}\rightarrow\mu^{+}+\nu$; or on Cerenkov-like emission of neutrino-antineutrino pairs, as in the normally forbidden process $p^{+}\rightarrow p^{+}+\nu+\bar{\nu}$, for which the kinematics are (because the neutrinos are nearly massless) essentially the same as those for a Cerenkov process.